# Lookup multivariate Kolmogorov-Arnold Networks

### v1

@@ Abstract
High-dimensional linear mappings, or linear layers, dominate both the parameter count and the computational cost of most modern deep-learning models. We introduce a general drop-in replacement, lookup multivariate Kolmogorov–Arnold Networks (lmKANs), which deliver a substantially better trade-off between capacity and inference cost. Our construction expresses a general high-dimensional mapping through trainable low-dimensional multivariate functions. These functions can carry dozens or hundreds of trainable parameters each, and yet it takes only a few multiplications to compute them because they are implemented as spline lookup tables. Empirically, lmKANs reduce inference FLOPs by up to 6.0× while matching the flexibility of MLPs in general high-dimensional function approximation. In another feedforward fully connected benchmark, on the tabular-like dataset of randomly displaced methane configurations, lmKANs enable more than 10× higher H100 throughput at equal accuracy. Within frameworks of Convolutional Neural Networks, lmKAN-based CNNs cut inference FLOPs at matched accuracy by 1.6–2.1× and by 1.7× on the CIFAR-10 and ImageNet-1k datasets, respectively. Our code, including dedicated CUDA kernels, is available online at https://github.com/schwallergroup/lmkan.
@@ 1 Introduction
With a sufficient amount of training data, the capabilities of deep-learning models systematically improve with the number of trainable parameters Zhai et al. (2022); Kaplan et al. (2020). However, deploying very large models is challenging because of the associated inference cost.
In most models, high-dimensional linear mappings dominate both the parameter count and the computational cost. Standard multilayer perceptrons (MLPs) alternate linear layers with activations, and sometimes with a few other layers Ioffe and Szegedy (2015); Hinton et al. (2012). If $N$ is the width of a layer, the parameter count and inference cost of these linear mappings scale as $\mathcal{O}(N^{2})$, whereas most other layers scale only as $\mathcal{O}(N)$. The same observation holds for many other architectures. Transformers Vaswani et al. (2017), when applied to very long sequences, are one of the few notable exceptions because the cost of attention grows quadratically with the number of tokens. Even in that case, however, the cost of the linear mappings remains substantial, not to mention the potential use of fast approximations of attention Choromanski et al. (2020).
The computational cost of a linear layer is proportional to the number of its parameters: at inference each parameter induces one multiplication per input, where input is a whole input object in the case of MLPs, token in the case of Recurrent Neural Networks Elman (1990) and Transformers Vaswani et al. (2017), node or edge in the case of Graph Neural Networks Zhou et al. (2020), patch of an image in case of Convolutional Neural Networks LeCun et al. (2002); Krizhevsky et al. (2012), and similarly for other architectures.
Spline lookup tables make it possible to do better than that. Consider, for example, a one-dimensional piecewise-linear function $f(x)$ on the interval from 0 to $1$ with a uniform grid. On each interval it is given as f(x)=a[i]x+b[i]f(x)=a[i]*x+b[i], where $i$ is the interval index. With $G$ intervals, the function has $2G$ parameters, out of which $G+1$ are independent once continuity at the internal grid points is enforced. Yet the computational cost of evaluating such a function at any given point is $\mathcal{O}(1)$, not depending on $G$. The computational pipeline involves first determining the current grid interval as $i=\lfloor x*G\rfloor$, and next evaluating only one linear function as f(x)=a[i]x+b[i]f(x)=a[i]*x+b[i].
Kolmogorov-Arnold Networks (KANs) Liu et al. (2024), designed as a general alternative to MLPs, are natural hosts for spline lookup tables as they construct a general high-dimensional mapping through a collection of trainable univariate functions.
@@
  • We propose lookup multivariate Kolmogorov-Arnold Networks (lmKANs) that are built upon multivariate low-dimensional functions organized as spline lookup tables for efficient inference. Ignoring a non-asymptotic $\mathcal{O}(N)$ term, the required inference FLOPs are only $2\times$ those of a linear layer of the same shape, while the number of trainable parameters can be up to several hundred times higher.
  • We implement custom CUDA kernels that enable efficient inference of lmKANs on modern GPUs. When using $8\times 8$ tile size, on the H100 GPU, our implementation enables up to $\approx 88\times$ better inference time per number of trainable parameters compared to that of a linear layer.
[…]
@@ 2 Related work
Kolmogorov-Arnold Representation Theorem (KART) Kolmogorov ; Arnold (2009) states that a continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as:
@@
where $\phi_{q,p}:[0,1]\to\mathbb{R}$, and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous univariate functions. It has been a long debate Girosi and Poggio (1989); Schmidt-Hieber (2021) on the usefulness of this theorem for machine learning because of the general non-smoothness and wild behavior of the inner functions. Nevertheless, it inspired the construction of Kolmogorov-Arnold Networks Hecht-Nielsen (1987); Igelnik and Parikh (2003), whose layers are defined as:
$y_{q}=\sum_{p}f_{qp}(x_{p})\text{,}$ (2)
where $f_{qp}$ are trainable functions. Liu et al. (2024) introduced the modern version, which suggests stacking an arbitrarily large number of KAN layers and using an arbitrarily large number of neurons, similarly to MLPs. While Liu et al. (2024) illustrates superior performance of KANs, it has a limitation that for many studied cases, KART or KART-like (matching KANs with more than one hidden layer and a larger number of neurons) representations for ground truth target are known in closed form, and are reasonably smooth. On the contrary, Yu et al. (2024) found that KANs often fall short compared to MLP for general tasks.
The idea of lookup-based $\mathcal{O}(1)$ computations of KAN univariate functions is sometimes briefly mentioned but rarely implemented in practice Somvanshi et al. (2024); Ji et al. (2024), likely because of challenges associated with an efficient GPU implementation. Surprisingly, most of the research goes in the somewhat opposite direction. B-splines, piecewise polynomial basis functions used in the original KAN paper, have compact support and thus are well suited for $\mathcal{O}(1)$ inference. Subsequent works often substitute them for dense basis functions, such as Chebyshev polynomials SS et al. (2024) or Fourier harmonics Xu et al. (2024). The case of FastKAN Li (2024a), which substitutes sparse B-splines with similarly-looking dense Gaussian Radial Basis functions exclusively for the sake of optimization, is especially notable.
A few works, such as Moradzadeh et al. (2024) and Huang et al. (2025), implement the lookup idea. Moradzadeh et al. (2024), however, predicts B-spline coefficients by MLP for the given grid points, which, thus, are not fully independent from each other. Additionally, it provides benchmarks against a naive implementation of KANs and not against MLPs, which are still state-of-the-art for general tasks. Huang et al. (2025) achieves remarkable efficiency on a small-scale problem from the original KAN paper by algorithm-hardware co-design using the TSMC 22 nm RRAM-ACIM chip.
In this work, we provide CUDA kernels for efficient inference and benchmark the introduced models against MLPs on general tasks where KART representations are not known and where there is no reason to believe that they are smoother than in other cases.
In Appendix A, we review further related work beyond KANs.
@@ 3 Lookup multivariate Kolmogorov-Arnold Networks
At first glance, given that the inference cost of spline lookup tables does not depend on the number of parameters, very expressive univariate functions with tens of thousands of trainable parameters each are an ideal match with the Kolmogorov-Arnold Representation Theorem. However, KANs rarely use more than a few dozen parameters per function in practice. The difference between a univariate function parameterized by tens of thousands of parameters and just a few dozen is the capability of the former to parametrize a very high frequency band. On the one hand, this expressivity is necessary for closely approximating the ’wild-behaviour’ of KART inner functions, but on the other, it raises concerns about training stability and generalization. On the contrary, multivariate functions can "accommodate" a significantly larger number of parameters without spilling expressive power to exceedingly high frequency bands. For instance, a four-dimensional function with just 10 grid points along each dimension has roughly the same number of trainable parameters as a univariate one with $\sim 10^{4}$ grid intervals.
@@
In Sec. 4.4 we empirically compare the two-dimensional version of lmKANs with one-dimensional FastKAN. The outcomes of these numerical experiments reinforce the intuitive considerations given here and suggest that multidimensional building blocks can indeed be more effective placeholders for a large number of parameters in a practical setup.
Additionally, it is worth noting that if necessary, multivariate functions $f_{qp}$ can always fall back to sums of univariate ones, which would make the whole lmKAN fall back to standard KAN. Thus, the Kolmogorov-Arnold Representation Theorem is applicable also to our construction.
@@ 3.1 Function parametrization
The one-dimensional sigma grid, which is illustrated in Fig. 3a, is generated by any sigmoid-like function $\sigma(x)$. If the desired number of grid intervals is $G$, then the grid points are given as the intersection of G1G-1 equispaced percentile levels with $\sigma(x)$. An example of piecewise linear function defined on such a grid is given in Fig. 3b. Such a construction spans the entire real axis. The grid has the finest resolution near the origin, and becomes progressively coarser as $|x|$ increases. For a given $x$, the index of the corresponding grid interval can be computed as $i=\lfloor\sigma(x)G\rfloor$, which makes such a grid suitable for $\mathcal{O}(1)$ computations.
@@
The choice of $\sigma(x)$ in the construction above is somewhat arbitrary. An additional consideration is that it would be beneficial to distribute inputs evenly, or approximately evenly, across the grid intervals. Consider, for instance, a pathological scenario when all the input $x$ values belong to a single grid interval. In such a case, the multitude of parameters that govern the function’s behavior on all the other intervals would never influence the output, leaving the parametrized function no more expressive than a two-parameter linear mapping.
If the probability distribution of inputs is known, then one way to ensure perfect balance is to set $\sigma(x)$ to be the corresponding cumulative distribution function (CDF). However, it is challenging to implement such a dynamic percentile grid in practice, as the distribution of inputs is unknown, can evolve during training, and the corresponding CDF can hardly be queried in $\mathcal{O}(1)$ time.
Instead, we keep activations in a controlled range. A batch-normalization layer without affine parameters placed before each lmKAN layer forces zero mean and unit variance. While it does not pin the entire distributions, it is safe to assume that they will be close enough to the standard normal for a reasonable ratio of neurons. Therefore, a viable solution would be to precede each lmKAN layer with a batch normalization with disabled affine transforms, and select $\sigma(x)$ to be the standard Gaussian CDF.
While the computational complexity of the standard Gaussian CDF is $\mathcal{O}(1)$, it appears to be quite slow from a non-asymptotic point of view. Thus, in order to ensure the utmost efficiency, we propose a simple approximation requiring the computation of only a single exponent function at inference, which is given in appendix B along with other details.
For multivariate functions we apply the same one-dimensional grid independently to each coordinate.
Figure 3: (a) Construction of the sigma grid; (b) an example of piecewise linear function defined on such a grid (c) Second-order B-spline (d) Two-dimensional second-order B-spline
@@
We use B-splines of second order built on top of the described grids as basis functions to parametrize the lmKAN inner functions. A one-dimensional second-order B-spline centered around grid point $i$ is given in Fig. 3c. It takes non-zero values on only two adjacent grid intervals around the center grid point. If there are $G$ intervals, then we use $G+1$ basis functions: G1G-1 such B-splines centered around each inner grid point, and two linear functions on the left-most and right-most infinite intervals. For any given point $x$ there are only two non-zero B-splines $B_{i}(x)$ - the ones with $i=\lfloor\sigma(x)G\rfloor$ and $i=\lfloor\sigma(x)G\rfloor+1$ (In corner case when $x$ coincide with one of the grid points, there is only one). Appendix B provides the definition of B-splines on edge intervals we use in this work along with other details.
@@
With such a construction, there are $(G+1)^{2}$ independent parameters per each 2D function, parametrized functions are bilinear on each 2D grid interval, all the functions are continuous for arbitrarily $p_{i_{1}i_{2}}$, and all but the edge coefficients $p_{i_{1}i_{2}}$ have a simple interpretation of a value of the function on the corresponding grid point.
For any given point $(x_{1},x_{2})$ there are only four non-zero B-splines; thus, one needs to evaluate only four terms of Eq. 4 to compute $f(x_{1},x_{2})$, which forms the basis of $\mathcal{O}(1)$ computations. The full algorithm to compute a standalone two-dimensional function is given in Appendix B.1.
@@ 3.2 Computational cost
Overall, the functional form of lmKANs involves computations of many low-dimensional functions with exactly the same arguments (those in the same column, see Fig. 2). When doing so, it is possible to reuse many of the intermediate values, such as indices of the grid intervals and B-splines. As Algorithm 1 elaborates, these intermediate values can be computed once for each pair of inputs, and then re-used to compute the value of a given function with just four multiply-add operations for the 2D case.
As the algorithm shows, the preamble part contributes only to an asymptotically insignificant $\mathcal{O}(N)$ term, where $N$ is the input, $N_{\text{in}}$, or the output, $N_{\text{out}}$, dimension. Given that the total number of 2D functions in an lmKAN layer is $N_{\text{in}}/2]N_{\text{out}}$, the total number of required multiply-add operations for the dominant, $\mathcal{O}(N^{2})$, part is $4[N_{\text{in}}/2]N_{\text{out}}=2N_{\text{in}}N_{\text{out}}$, just $\mathbf{2\times}$ that of a linear layer of the same shape. Following the common practice He et al. (2016), we estimate FLOPs as the number of fused multiply-adds of the main asymptotic term for both MLPs and lmKANs.

### v2

@@ Abstract
High-dimensional linear mappings, or linear layers, dominate both the parameter count and the computational cost of most modern deep-learning models. We introduce a general-purpose drop-in replacement, lookup multivariate Kolmogorov-Arnold Networks (lmKANs), which deliver a substantially better trade-off between capacity and inference cost. Our construction expresses a general high-dimensional mapping through trainable low-dimensional multivariate functions. These functions can carry dozens or hundreds of trainable parameters each, and yet it takes only a few multiplications to compute them because they are implemented as spline lookup tables. Empirically, lmKANs reduce inference FLOPs by up to 6.0× while matching the flexibility of MLPs in general high-dimensional function approximation. In another feedforward fully connected benchmark, on the tabular-like dataset of randomly displaced methane configurations, lmKANs enable more than 10× higher H100 throughput at equal accuracy. Within frameworks of Convolutional Neural Networks, lmKAN-based CNNs cut inference FLOPs at matched accuracy by 1.6–2.1× and by 1.7× on the CIFAR-10 and ImageNet-1k datasets, respectively. Our code, including dedicated CUDA kernels, is available online at https://github.com/schwallergroup/lmkan.
@@ 1 Introduction
With a sufficient amount of training data, the capabilities of deep-learning models systematically improve with the number of trainable parameters [1, 2]. However, deploying very large models is challenging because of the associated inference cost.
In most models, high-dimensional linear mappings dominate both the parameter count and the computational cost. Standard multilayer perceptrons (MLPs) alternate linear layers with activations, and sometimes with a few other layers [3, 4]. If $N$ is the width of a layer, the parameter count and inference cost of these linear mappings scale as $\mathcal{O}(N^{2})$, whereas most other layers scale only as $\mathcal{O}(N)$. The same observation holds for many other architectures. Transformers [5], when applied to very long sequences, are one of the few notable exceptions because the cost of attention grows quadratically with the number of tokens. Even in that case, however, the cost of the linear mappings remains substantial, not to mention the potential use of fast approximations of attention [6].
The computational cost of a linear layer is proportional to the number of its parameters: at inference, each parameter induces one multiplication per input, where the input is a whole input object in the case of MLPs, a token in the case of Recurrent Neural Networks [7] and Transformers [5], a node or an edge in the case of Graph Neural Networks [8], a patch of an image in the case of Convolutional Neural Networks [9, 10], and similarly for other architectures.
Spline lookup tables make it possible to do better than that. Consider, for example, a one-dimensional piecewise-linear function $f(x)$ on the interval from 0 to $1$ with a uniform grid. On each interval, it is given as f(x)=a[i]x+b[i]f(x)=a[i]*x+b[i], where $i$ is the interval index. With $G$ intervals, the function has $2G$ trainable parameters, out of which $G+1$ are independent once continuity at the internal grid points is enforced. Yet the computational cost of evaluating such a function at any given point is $\mathcal{O}(1)$, not depending on $G$. The computational pipeline involves first determining the current grid interval as $i=\lfloor x*G\rfloor$, and then evaluating only one linear function as f(x)=a[i]x+b[i]f(x)=a[i]*x+b[i].
Kolmogorov-Arnold Networks (KANs) [11], designed as a general alternative to MLPs, are natural hosts for spline lookup tables as they construct a general high-dimensional mapping through a collection of trainable univariate functions.
@@
  • We propose lookup multivariate Kolmogorov-Arnold Networks (lmKANs) that are built upon multivariate low-dimensional functions instead of the univariate ones that standard KANs employ. We empirically compare the 2D version of lmKANs with 1D FastKAN [12] and find that lmKANs are more accurate and easier to train.
  • We implement the inner functions as spline lookup tables. Ignoring a non-asymptotic $\mathcal{O}(N)$ term, the required inference FLOPs are only $2\times$ those of a linear layer of the same shape, while the number of trainable parameters can be dozens or hundreds of times higher.
  • We provide custom CUDA kernels that enable efficient inference of lmKANs on modern GPUs. When using the $8\times 8$ tile size, on the H100 GPU, our implementation enables up to $\sim 88\times$ faster inference than a linear layer with the same number of trainable parameters.
  • We empirically compare lmKANs and MLPs across diverse datasets, scales, and backbones, using varied experimental setups to obtain a comprehensive view of performance. Across these conditions, lmKANs are consistently Pareto-optimal with respect to inference FLOPs. The performance of lmKANs is summarized in Fig. 1.
@@ 2 Related work
Kolmogorov-Arnold Representation Theorem (KART) [13, 14] states that a continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as:
@@
where $\phi_{q,p}:[0,1]\to\mathbb{R}$, and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous univariate functions. There has been a long debate [15, 16] on the usefulness of this theorem for machine learning because of the general non-smoothness and wild behavior of the inner functions. Nevertheless, it inspired the construction of Kolmogorov-Arnold Networks [17, 18], whose layers are defined as $y_{q}=\sum_{p}f_{qp}(x_{p})$, where $f_{qp}$ are trainable functions. Liu et al. [11] introduced the modern version, which suggests stacking an arbitrarily large number of KAN layers and using an arbitrarily large number of neurons, similarly to MLPs. While Liu et al. [11] illustrated strong performance of KANs, many test cases involve ground-truth functions with known, reasonably smooth KART or KART-like (matching KANs with more than one hidden layer and a larger number of neurons) closed-form representations. Subsequent works such as Yang and Wang [19], Kundu et al. [20], and Kashefi [21] further reinforced the efficiency of KANs. On the contrary, Yu et al. [22] found that KANs can fall short compared to MLPs for some tasks.
The idea of lookup-based $\mathcal{O}(1)$ computations of KAN univariate functions is sometimes briefly mentioned but rarely implemented in practice [23, 24], likely because of challenges associated with an efficient GPU implementation. Surprisingly, most of the research goes in the somewhat opposite direction. B-splines, piecewise polynomial basis functions used in the original KAN paper, have compact support and thus are well suited for $\mathcal{O}(1)$ inference. Subsequent works often replace them with dense basis functions, such as Chebyshev polynomials [25] or Fourier harmonics [26]. The case of FastKAN [12], which replaces sparse B-splines with similar-looking dense Gaussian radial basis functions exclusively for the sake of optimization, is especially notable.
A few works, such as Moradzadeh et al. [27] and Huang et al. [28], implement the lookup idea. Moradzadeh et al. [27], however, predict B-spline coefficients using an MLP for the given grid points, which, thus, are not fully independent of each other. Huang et al. [28] achieve remarkable efficiency on a small-scale problem from the original KAN paper by algorithm-hardware co-design using the TSMC 22 nm RRAM-ACIM chip. Poluektov and Polar [29] and Polar and Poluektov [30] employ piecewise linear parametrization suitable for $\mathcal{O}(1)$ inference but do not focus on inference efficiency.
In this work, we provide CUDA kernels for efficient inference and benchmark the introduced models against MLPs on general tasks where KART representations are not known in closed form and where there is no reason to expect them to be smoother than in other cases.
@@ 3 Lookup multivariate Kolmogorov-Arnold Networks
At first glance, given that the inference cost of spline lookup tables does not depend on the number of parameters, very expressive univariate functions with tens of thousands of trainable parameters each are an ideal match for the Kolmogorov-Arnold Representation Theorem. However, KANs rarely use more than a few dozen parameters per function in practice. The difference between a univariate function parametrized by tens of thousands of parameters and just a few dozen is the capability of the former to parametrize a very high frequency band. On the one hand, this expressivity is necessary for closely approximating the ’wild behavior’ of KART inner functions, but on the other, it raises concerns about training stability and generalization. On the contrary, multivariate functions can "accommodate" a significantly larger number of parameters without spilling expressive power into exceedingly high frequency bands. For instance, a four-dimensional function with just 10 grid points along each dimension has roughly the same number of trainable parameters as a univariate one with $\sim 10^{4}$ grid intervals.
@@
In Sec. 4.4, we empirically compare the two-dimensional version of lmKANs with one-dimensional FastKAN. The outcomes of these numerical experiments reinforce the intuitive considerations given here and suggest that multidimensional building blocks can indeed be more effective hosts for a large number of parameters in a practical setup.
Additionally, it is worth noting that, if necessary, multivariate functions $f_{qp}$ can always fall back to sums of univariate ones, which would make the whole lmKAN fall back to standard KAN. Thus, the Kolmogorov-Arnold Representation Theorem is applicable also to our construction.
@@ 3.1 Function parametrization
The one-dimensional sigma grid, which is illustrated in Fig. 3a, is generated by any sigmoid-like function $\sigma(x)$. If the desired number of grid intervals is $G$, then the grid points are given as the intersection of G1G-1 equispaced percentile levels with $\sigma(x)$. An example of a piecewise linear function defined on such a grid is given in Fig. 3b. Such a construction spans the entire real axis. The grid has the finest resolution near the origin, and becomes progressively coarser as $|x|$ increases. For a given $x$, the index of the corresponding grid interval can be computed as $i=\lfloor\sigma(x)G\rfloor$, which makes such a grid suitable for $\mathcal{O}(1)$ computations.
@@
The choice of $\sigma(x)$ in the construction above is somewhat arbitrary. An additional consideration is that it would be beneficial to distribute inputs evenly, or approximately evenly, across the grid intervals. If the probability distribution of inputs is known, then one way to ensure perfect balance is to set $\sigma(x)$ to be the corresponding cumulative distribution function (CDF). However, it is challenging to implement such a dynamic percentile grid in practice, as the distribution of inputs is unknown, can evolve during training, and the corresponding CDF can hardly be queried in $\mathcal{O}(1)$ time.
Instead, we keep activations in a controlled range. A batch-normalization layer without affine parameters placed before each lmKAN layer forces zero mean and unit variance. While it does not pin the entire distributions, it is safe to assume that they will be close enough to the standard normal for a reasonable ratio of neurons. Therefore, a viable solution would be to precede each lmKAN layer with a batch normalization with disabled affine transforms, and select $\sigma(x)$ to be the standard Gaussian CDF. In practice, we implement a fast approximation that requires computing only a single exponent function at inference, see Appendix B. For multivariate functions, we apply the same one-dimensional grid independently to each coordinate.
Figure 3: (a) Construction of the sigma grid; (b) an example of a piecewise linear function defined on such a grid (c) Second-order B-spline (d) Two-dimensional second-order B-spline
@@
We use B-splines of second order built on top of the described grids as basis functions to parametrize the lmKAN inner functions. A one-dimensional second-order B-spline centered around grid point $i$ is given in Fig. 3c. It takes non-zero values on only two adjacent grid intervals around the center grid point. If there are $G$ intervals, then we use $G+1$ basis functions: G1G-1 such B-splines centered around each inner grid point, and two linear functions on the left-most and right-most infinite intervals. For any given point $x$, there are only two non-zero B-splines $B_{i}(x)$ - the ones with $i=\lfloor\sigma(x)G\rfloor$ and $i=\lfloor\sigma(x)G\rfloor+1$ (In the corner case when $x$ coincides with one of the grid points, there is only one). Appendix B provides the definition of B-splines on edge intervals we use in this work, along with other details.
@@
With such a construction, there are $(G+1)^{2}$ independent parameters for each 2D function, parametrized functions are bilinear on each 2D grid interval, all the functions are continuous for arbitrary $p_{i_{1}i_{2}}$, and all but the edge coefficients $p_{i_{1}i_{2}}$ have a simple interpretation as the value of the function on the corresponding grid point.
For any given point $(x_{1},x_{2})$, there are only four non-zero B-splines; thus, one needs to evaluate only four terms of Eq. 3 to compute $f(x_{1},x_{2})$, which forms the basis of $\mathcal{O}(1)$ computations. The full algorithm to compute a standalone two-dimensional function is given in Appendix B.1.
@@ 3.2 Computational cost
Overall, the functional form of lmKANs involves computations of many low-dimensional functions with exactly the same arguments (those in the same column, see Fig. 2). When doing so, it is possible to reuse many of the intermediate values, such as indices of the grid intervals and B-splines. As Algorithm 1 elaborates, these intermediate values can be computed once for each pair of inputs, and then utilized to compute the value of a given function with just four multiply-add operations for the 2D case.
As the algorithm shows, the preamble part contributes only to an asymptotically insignificant $\mathcal{O}(N)$ term, where $N$ is the input ($N_{\text{in}}$) or output ($N_{\text{out}}$) dimension. Given that the total number of 2D functions in an lmKAN layer is $N_{\text{in}}/2]N_{\text{out}}$, the total number of required multiply-add operations for the dominant, $\mathcal{O}(N^{2})$, part is $4[N_{\text{in}}/2]N_{\text{out}}=2N_{\text{in}}N_{\text{out}}$, just $\mathbf{2\times}$ that of a linear layer of the same shape. Following the common practice [31], we estimate FLOPs as the number of fused multiply-adds of the main asymptotic term for both MLPs and lmKANs.